Factorise the quadratic polynomial by splitting the middle term 
X^2-3x-5
 

Grade:10

2 Answers

ANIKET RAJ
5 years ago
you can do it by complete square method:
x2-3x-5
first check for coefficient of x2 whic is 1
x2+2*x*-3/2+(3/2)2-(3/2)2-5=0
(x-3/2)2 =(3/2)+5
(x-3/2)2=(√14/2)2
(x-3/2)2-(√14/2)2=0
a2-b2=(a+b)(a-b)
(x-3/2+√14/2)(x-3/2-√14/2) …................Ans
PLEASE APPROVE
Nivedita
5 years ago
This is another method of completing the square-
x2 - 3x -5
(x+ -3/2)2 – (-3/2)2 – 5 =0
(x+ -3/2)2 – (2.25) – 5 =0
(x+ -3/2)2 = 7.25
x + -3/2 = + - ∫7.25
x = +-∫7.25 – -3/2